# Characterizing the nuclear models informed by PREX and CREX: a view from Bayesian inference

### v1

@@ I.3 Weak Charge form factors with spin-orbit contributions
$b_{4}$ [$\mathrm{fm}^{4}$] [-0.36, 0.72]
$b^{\prime}_{4}$ [$\mathrm{fm}^{4}$] [-0.36, 0.72]
I.4 Prior
In this and the following two sections, we describe the prior, likelihood and Posterior in our Bayesian analysis. Our RMF model have 12 parameters. We fixed the $\delta$, $\omega$ and $\rho$ meson masses, $m_{\delta}=980$ MeV, $m_{\omega}=782.5$ MeV and $m_{\rho}=763$ MeV as in many FSU type models, leaving us with 9 free parameters. $\sigma$ meson mass is a free parameter with uniform distribution $[450,550]$ MeV. We also take uniform prior in SNM properties $n_{s}$, $BE$, $M^{*}$ and $K$, see Table 5. These 4 bulk nuclear properties can be mapped to 4 isoscalar coupling constant, as in Chen and Piekarewicz (2014), except for $\omega$ meson self-coupling $\zeta_{\omega}$ which take flat distribution $[0,0.03]$. Isovector sector is governed by $g_{\delta}$, $g_{\rho}$ and $\Lambda_{\omega\rho}$ which can be mapped to symmetry energy properties $S_{\mathrm{V}}$ and $L$, assuming a flat distribution of $g_{\delta}\in[0,1500]$. $S_{\mathrm{V}}$ take uniformly prior while $L$ is uniformly distributed in $[L^{-},L^{+}]$. Because $L^{-}$ and $L^{+}$ are not fixed but determined by other parameters. And flat distributions of these parameters finally result in a non-flat prior distribution of L,
In this and the following two sections, we describe the prior, likelihood, and posterior in our Bayesian analysis. Our RMF model has 12 parameters. We fixed the $\delta$, $\omega$, and $\rho$ meson masses at $m_{\delta}=980$ MeV, $m_{\omega}=782.5$ MeV, and $m_{\rho}=763$ MeV, as in many FSU-type models, leaving us with 9 free parameters. The $\sigma$ meson mass is a free parameter with a uniform distribution in the range $[450,550]$ MeV. We also take a uniform prior for the symmetric nuclear matter (SNM) properties $n_{s}$, $BE$, $M^{*}$, and $K$, as shown in Table 5. These 4 bulk nuclear properties can be mapped to 4 isoscalar coupling constants, as in Chen and Piekarewicz (2014), except for the $\omega$ meson self-coupling $\zeta_{\omega}$, which takes a flat distribution in the range $[0,0.03]$. The isovector sector is governed by $g_{\delta}$, $g_{\rho}$, and $\Lambda_{\omega\rho}$, which can be mapped to the symmetry energy properties $S_{\mathrm{V}}$ and $L$, assuming a flat distribution of $g_{\delta}$ in the range $[0,1500]$. $S_{\mathrm{V}}$ takes a uniform prior, while $L$ is uniformly distributed in $[L^{-},L^{+}]$. Because $L^{-}$ and $L^{+}$ are not fixed but determined by other parameters, the flat distributions of these parameters result in a non-flat prior distribution of $L$:
[…]
@@ I.4 Prior
The prior distribution of Skyrme parameters is chosen similarly to ensure consistency with the prior distribution used for the RMF models. We use flat prior distributions for $n_{s}$, $B$, $M^{*}$, $K$, and $S_{\mathrm{V}}$, as well as for the $x_{i}$, $b_{4}$, and $b^{\prime}_{4}$ parameters with bounds in Table 5. The range of the $x_{i}$, $b_{4}$, and $b_{4}^{\prime}$ parameters in the Skyrme prior distributions are determined as follows. First, we examined 11 widely used Skyrme models (SGII, NRAPR, UNEDF0, UNEDF2, SLy4, SV-min, SkO, SkOp, BSk16, KDE0v, and Gs). We then identified the maximum and minimum $x_{i}$, $b_{4}$, and $b^{\prime}_{4}$ parameters in these 11 models, denoted as $x_{i}^{\mathrm{max}}$, $x_{i}^{\mathrm{min}}$, $b_{4}^{\mathrm{max}}$, $b_{4}^{\mathrm{min}}$, $b_{4}^{\prime\mathrm{max}}$, and $b_{4}^{\prime\mathrm{min}}$. To cover a wide range of possible Skyrme parameters, we determine the prior distribution of $x_{i}$, $b_{4}$, and $b^{\prime}_{4}$ parameters as $[x_{i}^{\mathrm{min}}-\Delta x_{i},x_{i}^{\mathrm{max}}+\Delta x_{i}]$, $[b_{4}^{\mathrm{min}}-\Delta b_{4},b_{4}^{\mathrm{max}}+\Delta b_{4}]$, and $[b_{4}^{\prime\mathrm{min}}-\Delta b_{4}^{\prime},b_{4}^{\prime\mathrm{max}}+\Delta b_{4}^{\prime}]$, where $\Delta x_{i}=x_{i}^{\mathrm{max}}-x_{i}^{\mathrm{min}}$, $\Delta b_{4}=b_{4}^{\mathrm{max}}-b_{4}^{\mathrm{min}}$, and $\Delta b_{4}^{\prime}=b_{4}^{\prime\mathrm{max}}-b_{4}^{\prime\mathrm{min}}$.
@@
Table 6: The list of experiment and observation with adopted errors. Charge and weak form factor $F_{ch}$ and $F_{W}$ listed here correspond to momentum transfer $q=0.8733$ fm-1 for 48Ca in CREX or $0.3977$ fm-1 for in PREX.
@@
I.5 Likelihood
@@
In the Bayesian analysis of RMF and Skyrme models, we use the same constraints and uncertainties. To account for theoretical uncertainty, we incorporate a 2% uncertainty in the measured value of $R_{ch}$ and a 5% uncertainty in the standard deviation of $B/A$. It is also necessary to constrain $F_{ch}$ accurately to ensure that the electric weak form factor difference, $F_{ch}-F_{W}$, can be interpreted as a measurement in PREX and CREX. The constraints on $B/A$, $R_{ch}$, and $F_{ch}$ together form the basic nuclear constraints in our Bayesian analysis, as listed in Table 6.
@@
Sampling with basic nuclear constraints results in drastically different posteriors for $L$. In the Skyrme model, negative $L$ is suppressed by the constraint that the pressure of matter in $\beta$-equilibrium is always increasing with energy density. However, RMF models with negative $L$ around SNM can have a large symmetry energy slope around PNM due to the breakdown of the quadratic formula of symmetry energy with large $g_{\delta}$ and $g_{\rho}$. Since our goal is not to constrain the nuclear model with basic nuclear constraints, we add an additional likelihood factor to our RMF models to ensure the $S_{\mathrm{V}}$-$L$ distribution is compatible with only basic nuclei constraints, as shown by the black dashed ellipse in Fig. 1 in the letter.
In addition to the basic nuclei constraints, we impose additional PREX and CREX seperately and jointly, forming 4 sets of likelyhoods: basic nuclei, +PREX, +CREX and +PREX+CREX. We take the electric weak form factor difference $F_{ch}-F_{W}$ instead of weak form factor $F_{W}$ as constraints of PREX and CREX experiments as tabulated in Table 6. In principle, weak form factor $F_{W}$ can be determined from the parity violating asymmetry $A_{PV}$ measured in PREX and CREX experiments by the Coulomb distortion calculation assuming a electric charge distribution. As the electric charge distribution has already been determined experimentally with great accuracy, both $F_{W}$ and $F_{ch}-F_{W}$ are determined experimentally from the measured $A_{PV}$ without any model dependence. However, the theoretical models like RMF and Skyrme models optimized with charge radius doesn’t have enough accuracy to ensure consistency with the charge distribution assumed in Coulomb distortion calculation. Therefore, we need to make sure that the value of $F_{ch}$ used in our calculations is close to that used in analysis of PREX and CREX. To achieve this, we take $\sigma_{F_{ch}}=0.001$ which is much smaller than $\sigma_{F_{W}}$. And due to the remaining uncertainty in $F_{ch}$, we consider $F_{ch}-F_{W}$ in the likelihood for its better constraints than $F_{W}$.
@@
In addition to the basic nuclear constraints, we impose additional PREX and CREX constraints separately and jointly, forming four sets of likelihoods: basic nuclei, +PREX, +CREX, and +PREX+CREX. We take the electric weak form factor difference $F_{ch}-F_{W}$ instead of the weak form factor $F_{W}$ as constraints of PREX and CREX experiments, as tabulated in Table 6. In principle, the weak form factor $F_{W}$ can be determined from the parity-violating asymmetry $A_{PV}$ measured in PREX and CREX experiments through Coulomb distortion calculations, assuming an electric charge distribution. Since the electric charge distribution has already been determined experimentally with great accuracy, both $F_{W}$ and $F_{ch}-F_{W}$ are determined experimentally from the measured $A_{PV}$ without any model dependence. However, theoretical models like RMF and Skyrme models, optimized with charge radius, do not have enough accuracy to ensure consistency with the charge distribution assumed in Coulomb distortion calculations. Therefore, we need to make sure that the value of $F_{ch}$ used in our calculations is close to that used in the analysis of PREX and CREX. To achieve this, we take $\sigma_{F_{ch}}=0.001$, which is much smaller than $\sigma_{F_{W}}$. Due to the remaining uncertainty in $F_{ch}$, we consider $F_{ch}-F_{W}$ in the likelihood for its better constraints compared to $F_{W}$.
@@
Figure 4: The posterior distributions of model parameters, saturation properties and $\Delta F$ measured in PREX and CREX for Skyrme model (upper) and RMF (lower) model.
Figure 5: The Pearson correlation coefficient between model parameters, saturation properties and $\Delta F$ measured in PREX and CREX for Skyrme model (upper) and RMF (lower) model.
@@
I.6 Posterior
With the prior and likelihood discussed above, we obtain the posterior distributions of model parameters, saturation properties, and $\Delta F$ measured in PREX and CREX, as shown in 4. To highlight the significant aspects of the posterior distribution, we also plot the Pearson correlation coefficient between any two parameters, as shown in Fig. 5. Some model parameters, such as $t_{0}$, $t_{3}$, and $\alpha=\gamma-2$, are tightly correlated due to the mapping between saturation properties used in our prior. Isoscalar and isovector Yukawa couplings are also tightly correlated due to the priors on binding energy $B$ and effective mass $M^{*}$. For both Skyrme and RMF models, the symmetry energy $S_{\mathrm{V}}$ and its slope $L$ are very sensitive parameters related to the form factor difference measured in PREX and CREX. From the posterior of Bayesian analysis, we obtain the mean and covariance of $S_{\mathrm{V}}$ and $L$ with +PREX+CREX constraints,
@@ I.6 Posterior
Interestingly, despite the weak correlation between $b^{\prime}_{4}$ and $\Delta F^{\mathrm{Ca}48/\mathrm{Pb}208}$ in the Skyrme model, as observed in Reinhard et al. (2022b), $b^{\prime}_{4}$ can be quite sensitive to a linear combination of $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$, such as $a^{\prime\prime}\Delta F^{\mathrm{Ca}48}+b^{\prime\prime}\Delta F^{\mathrm{Pb}208}$, significantly influencing the distribution of $\Delta F^{\mathrm{Ca}48},\Delta F^{\mathrm{Pb}208}$. The posterior for $b^{\prime}_{4}$ increases to $b^{\prime}4=0.310^{+0.24}{-0.26}$ with additional PREX and CREX constraints from $b^{\prime}_{4}=0.244^{+0.21}{-0.19}$ with basic nuclei constraints only. A similar trend is also observed in RMF models. The detailed impact of $b^{\prime}_{4}$ on the nucleon distribution is discussed in a separate section below.
@@
[…]
Figure 9: Pearson covariance coefficient between neutron skin (form factor difference) of 48Ca (solid) and 208Pb (dashed), and pressure at various baryon number densities. Blue and Red correspond to RMF models and Skyrme models.
I.7 Linear Correlation Between $S_{\mathrm{V}}$, $L$, $\Delta F^{\mathrm{Ca48}}$, and $\Delta F^{\mathrm{Pb208}}$
@@
Figures 7-8 show the distributions of $(\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}},S_{\mathrm{V}})$ and $(\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}},L)$, where the linear correlations in Eqs. (5-6) in the letter are observed. The precision of the fitting formula in Eq. (5) remains within 10% (1-$\sigma$) for both Skyrme and RMF models, while Eq. (6) shows about a 10 MeV deviation for the Skyrme model and a 25 MeV deviation in the RMF model. An interesting trend is noted in Fig. 7 for both RMF and Skyrme models: the predicted points of ${\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}}}$ move collectively from left to right (towards the mean of joint PREX and CREX measurements) with increasing $S_{\mathrm{V}}$. This collective movement results from $S_{\mathrm{V}}$ being sensitive to the neutron skin of 208Pb while being insensitive to the neutron skin of 48Ca in both RMF and Skyrme models. The difference in sensitivity of $\Delta F^{\mathrm{Pb208}}$ and $\Delta F^{\mathrm{Ca48}}$ to $S_{\mathrm{V}}$ is reflected by the slope coefficients $a$ and $b$ in Eq. (5) in the letter, where $b$ is much larger than $a$ in both Skyrme and RMF fittings.

### v2

@@ Acknowledgements.
$b_{4}$ $\mathrm{fm}^{4}$] [-0.36, 0.72]
$b^{\prime}_{4}$ $\mathrm{fm}^{4}$] [-0.36, 0.72]
.4 Prior
In this and the following two sections, we describe the prior, likelihood, and posterior in our Bayesian analysis. Our RMF model has 12 parameters. We fixed the $\delta$, $\omega$, and $\rho$ meson masses at $m_{\delta}=980$ MeV, $m_{\omega}=782.5$ MeV, and $m_{\rho}=763$ MeV, as in many FSU-type models Chen and Piekarewicz (2014), leaving us with 9 free parameters. The $\sigma$ meson mass is a free parameter with a uniform distribution in the range $[450,550]$ MeV, corresponding to mass of $f_{0}(500)$Tanabashi et al. (2018). We also take a uniform prior for the symmetric nuclear matter (SNM) properties $n_{s}$, $B$, $M^{D}$, and $K$, as shown in Table 5. These 4 bulk nuclear properties can be mapped to 4 isoscalar coupling constants, as in Chen and Piekarewicz (2014), except for the $\omega$ meson self-coupling $\zeta_{\omega}$, which takes a flat distribution in the range $[0,0.03]$ , since $\zeta_{\omega}>0.03$ is suppressed by astronomical observations Huang et al. (2024), particularly due to the maximum-mass constraint. The isovector sector is governed by $g_{\delta}$, $g_{\rho}$, and $\Lambda_{\omega\rho}$, which can be mapped to the symmetry energy properties $S_{\mathrm{V}}$ and $L$, assuming a flat distribution of $g_{\delta}$ in the range $[0,1500]$, covering higtest value in DINO modelsReed et al. (2024). $S_{\mathrm{V}}$ takes a uniform prior, while $L$ is uniformly distributed in $L^{-},L^{+}]$. Because $L^{-}$ and $L^{+}$ are not fixed but determined by other parameters, the flat distributions of these parameters result in a non-flat prior distribution of $L$:
$P(L)=\int^{\zeta_{\omega}^{+}}_{\zeta_{\omega}^{-}}\dots\int^{m_{\sigma}^{+}}_{m_{\sigma}^{-}}\int_{L^{-}(m_{\sigma},\dots,\zeta_{\omega})}^{L^{+}(m_{\sigma},\dots,\zeta_{\omega})}\delta(L^{\prime}-L)dL^{\prime}dm_{\sigma}\dots d\zeta_{\omega}$ (36)
@@
The prior distribution of Skyrme parameters is chosen similarly to ensure consistency with the prior distribution used for the RMF models. We use flat prior distributions for $n_{s}$, $B$, $m^{*}$, $K$, and $S_{\mathrm{V}}$, as well as for the $x_{i}$, $b_{4}$, and $b^{\prime}_{4}$ parameters with bounds in Table 5. The range of the $x_{i}$, $b_{4}$, and $b_{4}^{\prime}$ parameters in the Skyrme prior distributions are determined as follows. First, we examined 11 widely used Skyrme models (SGIIvan Giai and Sagawa (1981), NRAPRSteiner et al. (2005b), UNEDF0Kortelainen et al. (2014), UNEDF2Kortelainen et al. (2010b), SLy4Chabanat et al. (1998b), SV-minKlüpfel et al. (2009), SkOReinhard et al. (1999), SkOpDanielewicz and Lee (2009), BSk16Chamel et al. (2008), KDE0vAgrawal et al. (2005), and GsFriedrich and Reinhard (1986)). We then identified the maximum and minimum $x_{i}$, $b_{4}$, and $b^{\prime}_{4}$ parameters in these 11 models, denoted as $x_{i}^{\mathrm{max}}$, $x_{i}^{\mathrm{min}}$, $b_{4}^{\mathrm{max}}$, $b_{4}^{\mathrm{min}}$, $b_{4}^{\prime\mathrm{max}}$, and $b_{4}^{\prime\mathrm{min}}$. To cover a wide range of possible Skyrme parameters, we determine the prior distribution of $x_{i}$, $b_{4}$, and $b^{\prime}_{4}$ parameters as $x_{i}^{\mathrm{min}}-\Delta x_{i},x_{i}^{\mathrm{max}}+\Delta x_{i}]$, $b_{4}^{\mathrm{min}}-\Delta b_{4},b_{4}^{\mathrm{max}}+\Delta b_{4}]$, and $b_{4}^{\prime\mathrm{min}}-\Delta b_{4}^{\prime},b_{4}^{\prime\mathrm{max}}+\Delta b_{4}^{\prime}]$, where $\Delta x_{i}=x_{i}^{\mathrm{max}}-x_{i}^{\mathrm{min}}$, $\Delta b_{4}=b_{4}^{\mathrm{max}}-b_{4}^{\mathrm{min}}$, and $\Delta b_{4}^{\prime}=b_{4}^{\prime\mathrm{max}}-b_{4}^{\prime\mathrm{min}}$.
@@
Table 6: The list of experiment and observation with adopted errors. Mean value of the binding energy per nucleonWang et al. (2012) and charge radiiAngeli and Marinova (2013). Charge and weak form factor $F_{ch}$ and $F_{W}$ listed here correspond to momentum transfer $q=0.8733$ fm-1 for 48Ca in CREX Adhikari et al. (2022b) or $0.3977$ fm-1 for in PREX Adhikari et al. (2021b).
@@
.5 Likelihood
Figure 4: The likelihood posterior samples of Skyrme and RMF models colored by the parameter $b^{\prime}_{4}$ from -0.36 (deep blue) to 0.9 (yellow) $\mathrm{fm^{4}}$. The x-axis shows the log-likelihood from basic nuclei constraints including binding energy, charge radii, and charge form factor, while the y-axis shows the log-likelihood of PREX-2/CREX experiments in Table 6.
@@
In the Bayesian analysis of RMF and Skyrme models, we use the same constraints and uncertainties. To account for theoretical uncertainty, we incorporate a 2% uncertainty in the measured value of $R_{ch}$ and a 5% uncertainty in the standard deviation of $B/A$. It is also necessary to constrain $F_{ch}$ accurately to ensure that the electric weak form factor difference, $F_{ch}-F_{W}$, can be interpreted as a measurement in PREX-2/CREX. The constraints on $B/A$, $R_{ch}$, and $F_{ch}$ together form the basic nuclear constraints in our Bayesian analysis, as listed in Table 6.
@@
Sampling with basic nuclear constraints results in drastically different posteriors for $L$. In the Skyrme model, negative $L$ is suppressed by the constraint that the pressure of matter in $\beta$-equilibrium is always increasing with energy density. However, RMF models with negative $L$ around SNM can have a large symmetry energy slope around PNM due to the breakdown of the quadratic formula of symmetry energy with large $g_{\delta}$ and $g_{\rho}$. Since our goal is not to constrain the nuclear model with basic nuclear constraints, we add a likelihood factor to our RMF models to ensure the $S_{\mathrm{V}}$-$L$ distribution is compatible between Skyrme and RMF models with only basic nuclei constraints, as shown by the horizontally shaded ellipse in Fig. 1 in the letter.
@@
In addition to the basic nuclear constraints, we impose additional PREX-2/CREX constraints separately and jointly, forming four sets of likelihoods: basic nuclei, +PREX, +CREX, and +PREX+CREX. We take the electric weak form factor difference $F_{ch}-F_{W}$ instead of the weak form factor $F_{W}$ as constraints of PREX-2/CREX experiments, as tabulated in Table 6. In principle, the weak form factor $F_{W}$ can be determined from the parity-violating asymmetry $A_{PV}$ measured in PREX-2/CREX experiments through Coulomb distortion calculations, assuming an electric charge distribution. Since the electric charge distribution has already been determined experimentally with great accuracy, both $F_{W}$ and $F_{ch}-F_{W}$ are determined experimentally from the measured $A_{PV}$ without any model dependence. However, theoretical models like RMF and Skyrme models, optimized with charge radius, do not have enough accuracy to ensure consistency with the charge distribution assumed in Coulomb distortion calculations. Therefore, we need to make sure that the value of $F_{ch}$ used in our calculations is close to that used in the analysis of PREX-2/CREX. To achieve this, we take $\sigma_{F_{ch}}=0.005$, which is much smaller than $\sigma_{F_{W}}$. Due to the remaining uncertainty in $F_{ch}$, we consider $F_{ch}-F_{W}$ in the likelihood of its better constraints compared to $F_{W}$.
Figure 4 shows the likelihood of basic nuclei constraints vs the likelihood of PREX-2/CREX experiments listed in Table. 6. The lack of points in the upper right corner indicates the tension between basic nuclei properties and the PREX-2/CREX results. The fact that the points close to the upper right corner all have relatively large $\mathrm{b^{\prime}_{4}}$ indicates that larger $\mathrm{b^{\prime}_{4}}$ in mean field models may mitigate the tension between basic nuclei constraints and the PREX-2/CREX experiments.
@@
[…]
Figure 5: The posterior distributions of model parameters, saturation properties and $\Delta F$ measured in PREX-2/CREX for Skyrme model. The Pearson correlation coefficient between each pair of variables is displayed in the upper right corner.
[…]
Figure 6: The same as Fig.5 but for RMF model.
@@
.6 Posterior
With the prior and likelihood discussed above, we obtain the posterior distributions of model parameters, saturation properties, and $\Delta F$ measured in PREX-2/CREX, as shown in Fig. 5 and Fig. 6. To highlight the significant aspects of the posterior distribution, we also plot the Pearson correlation coefficient between any two parameters, as shown in the upper right corner of in Fig. 5 and Fig. 6. Some model parameters, such as $t_{0}$, $t_{3}$, and $\alpha=\gamma-2$, are tightly correlated due to the mapping between saturation properties used in our prior. Isoscalar and isovector Yukawa couplings are also tightly correlated due to the priors on binding energy $B$ and effective mass $M^{*}$. For both Skyrme and RMF models, the symmetry energy $S_{\mathrm{V}}$ and its slope $L$ are very sensitive parameters related to the form factor difference measured in PREX-2/CREX. From the posterior of Bayesian analysis, we obtain the mean and covariance of $S_{\mathrm{V}}$ and $L$ with +PREX+CREX constraints,
@@
Interestingly, despite the weak correlation between $b^{\prime}_{4}$ and $\Delta F^{\mathrm{Ca}48/\mathrm{Pb}208}$ in the Skyrme model, as observed in Reinhard et al. (2022b), $b^{\prime}_{4}$ can be quite sensitive to a linear combination of $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$, such as $a^{\prime\prime}\Delta F^{\mathrm{Ca}48}+b^{\prime\prime}\Delta F^{\mathrm{Pb}208}$, significantly influencing the distribution of $\Delta F^{\mathrm{Ca}48},\Delta F^{\mathrm{Pb}208}$. The posterior for $b^{\prime}_{4}$ increases to $b^{\prime}_{4}=0.310^{+0.24}_{-0.26}$ with additional PREX-2/CREX constraints from $b^{\prime}_{4}=0.244^{+0.21}_{-0.19}$ with basic nuclei constraints only. A similar trend is also observed in RMF models. The detailed impact of $b^{\prime}_{4}$ on the nucleon distribution is discussed in a separate section below.
@@
We don’t expect isoscalar parameters such as binding energy $B$ incompressibility $K$, self coupling of $\omega$ meson and $\zeta$ being sensitive to PREX and CREX which probes isovector properties. Therefore, their posterior is relatively flat as in the prior. $\Lambda_{\omega\rho}$ represents meson mixing between the two vector mesons. Therefore, it’s extremely correlated with the Yukawa coupling of the two vector mesons. The meson mixing is small when Yukawa couplings of the two vector mesons are large.
.7 Confidence level
The aim of this section is to quantify the level of consistency between theoretical predictions of $\Delta F$ for 48Ca and 208Pb, derived from Skyrme and RMF models, and the experimental measurements reported by PREX-2/CREX. This is accomplished using the well-established statistical framework of confidence levels, which quantify the degree to which new measurements are compatible with existing knowledge.
Let us assume that the state of knowledge prior to the CREX measurement suggests $\Delta F^{{}^{48}\mathrm{Ca}}=0.0432\pm 0.01$. Under this assumption, we expect that approximately 68% of subsequent measurements will fall within the 1-$\sigma$ confidence interval: $\Delta F^{{}^{48}\mathrm{Ca}}\in[0.0332,0.0532]$, assuming the knowledge is correct. Suppose an experiment yields $\Delta F^{{}^{48}\mathrm{Ca}}=0.0332$, which lies exactly at the lower bound of the 1-$\sigma$ interval. In that case, the probability that the current theoretical knowledge is correct, given the measurement, is 68%, while there is a 32% probability that the result is a statistical fluctuation inconsistent with the model—this defines the 1-$\sigma$ confidence level. Now, consider CREX measurement with error bar yields $\Delta F^{{}^{48}\mathrm{Ca}}=0.0277\pm 0.0055$. This result lies outside the original 1-$\sigma$ bound, and its 1-$\sigma$ upper bound barely overlaps with the 1-$\sigma$ lower bound of the theoretical expectation. It is a common misconception that such a scenario is equivalent to the prior one. In deed, the right approach involves computing the difference between the experimental result and the theoretical prediction as a probability distribution: $\mathcal{N}(\mu=0.0432-0.0277,\sigma=\sqrt{0.01^{2}+0.0055^{2}})$. The null hypothesis (i.e., that the discrepancy is due to statistical fluctuation) corresponds to zero deviation. In this case, zero lies approximately 1.36 standard deviations away from the mean of the difference distribution, implying a p-value (significance) of 17.6%. In other words, there is an 82.4% confidence level that the measurement is inconsistent with prior knowledge, once uncertainties are fully accounted for.
The previous example is straightforward because both knowledge and measurement are modeled as Gaussian distributions, and only a single observable is involved. For more general cases—arbitrary distributions with n-dimensional observables $\vec{x}$, we need to perform an overlapping integral to determine how confident level of our knowledge $P_{1}(\vec{x})$ is correct given the measurement $P_{2}(\vec{x})$. Bayesian confidence level can be defined as,
$P(\Delta\vec{x})$ $=$ $\int P_{2}(\vec{x})P_{1}(\vec{x}+\Delta\vec{x})d^{n}\vec{x}$ (40)
$1-\alpha$ $=$ $\frac{\int_{P(\Delta\vec{x})>P(0)}P(\Delta\vec{x})d^{n}\Delta\vec{x}}{\int P(\Delta\vec{x})d^{n}\Delta\vec{x}}$ (41)
where $1-\alpha$ quantifies the degree of incompatibility between knowledge and experiment. In the special case where $P_{2}$ is a delta function centered at $\vec{x}_{0}$, the overlap reduces to:
$P(\Delta\vec{x})$ $=$ $\int\delta(\vec{x}-\vec{x_{0}})P_{2}(\vec{x}+\Delta\vec{x})d^{n}\vec{x}=P_{2}(\vec{x_{0}}+\Delta\vec{x})$
then $\alpha(\vec{x}_{0})$ reduce to the significance level defined as the tail probability of highest posterior density region of $P_{2}(\vec{x})<P_{2}(\vec{x}_{0})$ as disscuss in the previous example with Gaussian distribution. Note that this analysis applies to probability distributions $P_{1}$ and $P_{2}$ that are localized and exhibit vanishing tails. The numerical implementation can be found in the GithubZhao .
Table 7 presents the resulting confidence levels $1-\alpha$ quantifying the agreement between theoretical predictions of $\Delta F$ from Skyrme and RMF models and the PREX-2/CREX measurements. The theoretical predictions are taken from the posterior distributions discussed in the previous section.
Overall, we observe greater tension between Skyrme predictions and experimental data compared to RMF predictions. This discrepancy primarily arises from the inclusion of the scalar-isovector $\delta$ meson in RMF models, with its large Yukawa coupling ($g_{\delta}^{2}\in[0,1500]$) enhancing the isovector spin-orbit interaction. This leads to an effective spin-orbit parameter $b_{4}^{\prime}\sim 1$ fm4, whereas the prior used in Skyrme modeling, $b_{4}^{\prime}\in[-0.36,0.72]$ fm4, though broader than typical ranges, remains insufficient to reproduce the central values of PREX-2/CREX.
Furthermore, the overlaps between knowledge and experiment are notably higher when only PREX is included, and lower when CREX or both measurements are considered. This indicates that the PREX-2 data is broadly consistent with current theoretical priors and imposes minimal update on the posterior, while the CREX measurement is in strong tension with prior expectations.
Table 7: Confidence levels ($1-\alpha$) quantifying the compatibility between posteriors of $\Delta F$ for 48Ca and 208Pb based on Skyrme and RMF models and the PREX-2/CREX measurements. Higher values correspond to increased statistical tension between model and experiment.
Posterior RMF Skyrme
basic nuclei 97.0% 99.7%
+PREX 97.4% 99.3%
+CREX 75.8% 97.6%
+PREX+CREX 69.7% 93.5%
@@
Figure 9: Similar to Fig. 3 in the letter but colored by symmetry energy parameterLL. Colorful lines corresponds to the minimum $\chi^{2}$ fitting of Eq. (6) in the letter with $L=[0,25,50,75,100,125,150]$ MeV.
Figure 10: Similar to Fig. 3 in the letter but without colored posterior points for RMF and Skyrme models.
Figure 11: Pearson correlation coefficient between neutron skin (form factor difference) of 48Ca (solid) and 208Pb (dashed), and pressure (upper panel) or symmetry energy slope (lower panel) at various baryon number densities. Blue and Red correspond to RMF models and Skyrme models.
.8 Linear Correlation Between $S_{\mathrm{V}}$, $L$, $b^{\prime}_{4}$, $\Delta F^{\mathrm{Ca48}}$, and $\Delta F^{\mathrm{Pb208}}$
@@
Figures 8-9 show the distributions of $(\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}},S_{\mathrm{V}})$ and $(\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}},L)$, where the linear correlations in Eqs. (5-6) in the letter are observed. The precision of the fitting formula in Eq. (5) remains within 10% (1-$\sigma$) for both Skyrme and RMF models, while Eq. (6) shows about a 10 MeV deviation for the Skyrme model and a 25 MeV deviation in the RMF model. An interesting trend is noted in Fig. 8 for both RMF and Skyrme models: the predicted points of $\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}}$} move collectively from left to right (towards the mean of joint PREX-2/CREX measurements) with increasing $S_{\mathrm{V}}$. This collective movement results from $S_{\mathrm{V}}$ being sensitive to the neutron skin of 208Pb while being insensitive to the neutron skin of 48Ca in both RMF and Skyrme models. The difference in sensitivity of $\Delta F^{\mathrm{Pb208}}$ and $\Delta F^{\mathrm{Ca48}}$ to $S_{\mathrm{V}}$ is reflected by the slope coefficients $a$ and $b$ in Eq. (5) in the letter, where $b$ is much larger than $a$ in both Skyrme and RMF fittings.